Ammonia gas is passed into water yielding a solution of density 0.93g/cm^3 and containing 18.6% ammonia by weight. The mass of ammonia per cc of solution is

Grade:11

5 Answers

Vikas TU
6 years ago
Dear Student,
We realize that thickness = mass/volume 
Accordingly, mass = thickness * volume 
Given thickness = 0.93g/cc 
Give us a chance to ascertain mass of 1 cc of arrangement 
mass = 0.93 * 1 = 0.93g 
Presently, we realize that 1 cc of arrangement weighs 0.93g 
Given that the arrangement contains 18.6% smelling salts by weight, 1 cc of arrangement ought to have: 
18.6% of 0.93g as smelling salts 
= 0.186 * 0.93 = 0.173g 
In this manner, the mass of Ammonia per cc of arrangement = 0.173g.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Himanshu Malik
6 years ago
We know that density = mass/volumeTherefore, mass = density * volumeGiven density = 0.93g/ccLet us calculate mass of 1 cc of solutionmass = 0.93 * 1 = 0.93gNow, we know that 1 cc of solution weighs 0.93gGiven that the solution contains 18.6% ammonia by weight, 1 cc of solution should have:18.6% of 0.93g as ammonia= 0.186 * 0.93 = 0.173gTherefore, the mass of Ammonia per cc of solution = 0.173g.
Shubham
6 years ago
18.6% NH3 in 100g solution density=0.93Now density= mass/volume Volume=density /mass= 100/0.93100/0.93 -----------------------18.61----------------------------------18.6/100/0.93=0.1729g/cm^3
Om singh
5 years ago
Mass percent of ammonia=18.6%
So, 18.6=(Mass of ammonia/mass of solution) *100
18.6=(mass of ammonia/volume of soln*density)*100
Mass of ammonia/volumeof soln=18.6*density*100=18.6*0.93*100=0.173g/ml.
The mass of NH3 is 0.173g/ml.
Hope you. 
Dont need to realize, only use siple formula. 
 
mysoreammonia mysoreammonia
5 years ago
Mumbai Ammonia & Chemical company is a major producer of ammonia solution and distributor or exporters of ammonia from India.